OpenStudy (anonymous):

The area of a triangular sheet of paper is 69 square inches. One side of the triangle is 5 inches more than three times the length of the altitude to that side. Find the length of that side and the altitude to the side.

OpenStudy (dtan5457):

woops give me a sec i made a mistake

OpenStudy (dtan5457):

let the altitude be x and the height be 3x+5 69=(3x+5)(x)(1/2) 138=3x^2+5x 3x^2+5x-138=0 solve that trinomial

OpenStudy (dtan5457):

do you know how to use the quadratic formula? @Ehoyle26

OpenStudy (dtan5457):

gonna need a response for further help :)

OpenStudy (anonymous):

no, I'm barely getting back into math again after almost 10 years

OpenStudy (dtan5457):

quadratic formula= \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a}\]

OpenStudy (dtan5457):

a=3 b=5 c=-138 in your case

OpenStudy (dtan5457):

\[\frac{ -5\pm \sqrt{25+1656} }{ 6 }\]

OpenStudy (dtan5457):

solve that

OpenStudy (anonymous):

So x= -23/3 and x=6

OpenStudy (dtan5457):

you will discard the negative part as it's not possible to have a negative length of a triangle

OpenStudy (dtan5457):

x=6 let the altitude be x and the height be 3x+5 plug x=6 back into to get altitude and height

OpenStudy (anonymous):

so, the altitude is 6 and the height is 23?

OpenStudy (dtan5457):

Yes

OpenStudy (anonymous):

Thank you so much! I appreciate your help!